OpenStudy (shadyykatie):

A stop sign is in the shape of a regular octagon. What is the measure of each angle? A. 115° B. 170° C. 135° D. 120° 2. Find the value of x. A. 10 B. 8.3 C. 7 D. 5 3. Find the measure of angle 1. A. 70 B. 35 C. 55 D. 60

OpenStudy (shadyykatie):

# 3

OpenStudy (shadyykatie):

# 2

OpenStudy (shadyykatie):

@arabpride

OpenStudy (arabpride):

for #3 110 + (1) = 180 -110 -110 (1) = 70 degreez ~~~~~~~~~~~~ for #2 20 = 2(3x - 5) 20 = 6x - 10 +10 +10 30 = 6x /6 /6 x = 5

OpenStudy (anonymous):

#1 Evaluate the following where n = 8 180 ((n - 2)/n)

OpenStudy (arabpride):

and since an octagon'z angle measure add up to 1080, you have to divide 1080 by 8 which givez you 135, so your answer is 135 degreez

OpenStudy (arabpride):

@ShadyyKatie

OpenStudy (shadyykatie):

im here! thank y'all so much I have more questions but I have to finish them by this afternoon

OpenStudy (arabpride):

that's fine you can just post them individually, tag me and ill help ya!! *^.^*

OpenStudy (shadyykatie):

Than You so much ur a lifesaver! <--- in a good way

OpenStudy (arabpride):

NOO Problem!! Glad to help ya out *^.^*